OpenStudy (anonymous):

f(x)=x^2 -1 / x^3 Determine the concavity of the function's curve, and find the inflection points.

OpenStudy (anonymous):

did you find the second derivative?

OpenStudy (anonymous):

no just first derivative

OpenStudy (anonymous):

if it was me, i would start with \[f(x)=\frac{1}{x}-\frac{1}{x^3}\] and find the derivative twice

OpenStudy (anonymous):

the first derivative was 

OpenStudy (anonymous):

what is second ?

OpenStudy (anonymous):

like i said, since you have to find the second derivative, i would start with \[f(x)=\frac{1}{x}-\frac{1}{x^3}\] so \[f'(x)=-\frac{1}{x^2}+\frac{3}{x^4}\] making the second derivative \[f''(x)=\frac{2}{x^3}-\frac{12}{x^5}\] snappier that way

OpenStudy (anonymous):

course now you still have to subtract, giving \[f''(x)=\frac{2x^2-12}{x^5}=\frac{2(x^2-6)}{x^5}\] and now the zeros are easy to find

OpenStudy (anonymous):

ok when I find second derivative what next ?

OpenStudy (anonymous):

find the critical points, where it is equal to zero or undefined. it is undefined at \(x=0\) but you can ignore that because the function is also undefined there, so it would be a miracle if the derivative existed at 0 the zeros of \(x^2-6\) you find in your head

OpenStudy (anonymous):

can you draw the solution ?

OpenStudy (anonymous):

not sure what you mean the zeros i.e. the inflection points are at \(x=-\sqrt6\) and \(x=\sqrt6\)

OpenStudy (anonymous):

aha ok than ?

OpenStudy (anonymous):

what else is there to do? you were asked for the inflection points right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks a lot :):)

OpenStudy (anonymous):

yw oh if you also need concavity, check for the signs of the second derivative, where is it positive or negative that is all

OpenStudy (anonymous):

ok :) thanks agine